how to solve mole concept quastion? My mole concept is not cleare. Is you help me to solve my this problem

Grade:11

2 Answers

mushtaq ahmed
6 years ago
Mole is nothing it is a unit in which we measure atoms just like dozen For banana its value is 6.62*10^32 which is avagrado`s number
Chhavi Jain
6 years ago
Dear student , 
 
Mole concept is a very basic topic of chemistry. It should be clear enough to even solve the questions of other related chapters.
 
Mole is the counting unit of chemistry .
 1 mole of any substance  = 6.022 * 1023  species or entities   
             = This is known as Avogadro’s number or Avogadro’s constant  ( denoted as NA  or  No )
Note :  these species can be atoms or  molecules or ions or electrons or protons , etc. 
 
Now question arises from where we get this value of 1 mole ?
Definition of 1 mole :   1 mole is defined as the no. of atoms present in 12 gram of C-12 isotope.
Let us take some examples :
(i)  1 mole of Mg atoms6.022 * 1023  Mg atoms (  you can imagine it is such a large no. of atoms of Mg, so it is rather easy to say 1 mole of Mg )
(ii)  5 moles of Mg atoms = 5 * 6.022 * 1023   Mg atoms 
(iii)  1 mole of CO2  molecules  =  6.022 * 1023  or NA   molecules of CO2
(iv)  4 moles  of   NH4+  ions =  4NA  ions of NH4+
 
Hope it is clear now. Still if you have further queries you can ask.
Chemistry Consultant
Askiitians